Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-02-23
ISS On-Orbit Status 02/23/08

Saturday -- off-duty day for CDR Whitson, FE-1 Malenchenko and FE-2 Eyharts except for housekeeping and voluntary work.

The crew performed the regular weekly three-hour task of thorough station cleaning. ("Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the CDR's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.)

CDR Whitson prepared the MELFI (Minus Eighty Degree Laboratory Freezer for the ISS) for the next DCB (Double Cold Bag) sample stow activity, first swapping box modules in Dewar 3 trays A & B, then inserting 16 -32deg ICEPAC belts in Dewar 3 (The MELFI dewar is currently powered off, i.e., at ambient temperature, so Peggy did not have to hurry through the normal 60 sec time 'window' after its opening.)

The CDR reconfigured the auto-rebooted PCS (Portable Computer System) laptops in COL (Columbus Orbital Laboratory), Lab and SM for use after the new R5 software patch uplink by the ground. Later in the day, she powered down the currently not needed Cupola and Airlock PCS laptops.

Peggy also finished up her assistance to ground specialists in troubleshooting the CDRA (Carbon Dioxide Removal Assembly) which had exhibited an unexpected differential pressure (dP) offset between half cycle 1 and half cycle 2 operations. (The troubleshooting, begun yesterday, involves inspecting and cleaning CDRA locations that could contribute to the dP anomaly at five locations that are accessible from the front of the rack and one area that requires removal of the TCCS (Trace Contaminant Control System) and use of the ULD (Ultrasonic Leak Detector). Yesterday Whitson experienced difficulties with the onboard fiber scope tool required for inspecting the AD/DES #2 (Absorbent/Desorbent Bed #2), resulting in deferral of about one hour of troubleshooting to today. Ground controllers plan to activate CDRA after dP troubleshooting is complete.)

The FE-1 completed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM.

Also, working from the Russian voluntary 'time permitting' task list, Yuri Malenchenko --

Completed another radiation data monitoring & logging session for flow & dose power data with the Matryoshka-R radiation payload and its LULIN-5 electronics box (accumulated readings were recorded on a log sheet for subsequent downlink to TsUP/Moscow via the BSR-TM payload data channel),
Conducted the regular checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at +20 degC, including a temperature check on its ART (automatic temperature recorder) (this is a daily monitoring/temp checking, carried on the Russian voluntary task list for the duration of Expedition 16),
Performed a session of the Russian "Uragan" (hurricane) earth-imaging program, using the Nikon D2X digital camera with 800 mm focal length lens (targets uplinked for today were the Western slopes of Carpathian mountains, Aral Sea and a dust storm (if present), the Kerch Strait area, the Main Caucasus mountain range (Elbrus volcano, Kazbek volcano), oil slicks in the Caspian Sea, the Pyrenees mountains( series of snapshots along the ridge)), and
Completed another KPT-3 session to make observations and take aerial KPT-3 photography of environmental conditions for Russia's Environmental Safety Agency (ECON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens.
Two new jobs added to Peggy Whitson's voluntary 'job jar' task list concern preparations for 1J/A EVAs, viz., (1) configuring some foam brushes and the tile sample bag, and (2) EMU resizing for 1J/A (also hard-scheduled next Wednesday).

At ~5:55am EST, Peggy & Leo held the (now) regular weekly audio tagup with ESA's Col-CC (Columbus Control Center) in Oberpfaffenhofen near Munich/Germany. (S/G-2 (Space-to-Ground 2) phone patch via SSC-10 (Station Support Computer 10)).

At ~7:40am, Yuri Malenchenko had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop).

At ~9:15am, the crew held its regular weekly planning conference (WPC) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

MSS Software Transition Update: Last night, ground controllers successfully completed uplink and checkout of the Robotics MSS (Mobile Service System) R5 software to the Prime C&C MDM (Command & Control Multiplexer/Demultiplexer computer). (This software update is required to incorporate Dextre/SPDM (Special Purpose Dexterous Manipulator) functionality into the MSS. In addition, there were a few updates to the SSRMS (Space Station Remote Manipulator System) and MBS (Mobile Base System) behavior as well as new SPDM PCS displays. The crew's next Robotics operations will occur on the new MSS R5 software. In prep for that, the ground will uplink a brief summary of system behavior changes for the DOUG (Dynamic Onboard Ubiquitous Graphics) software review for those ops.)

Columbus Update: Columbus systems continue to operate nominally.

Weekly Science Update (Expedition Sixteen -- Week 18)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Radiation measurements continue to be performed in the PIRS module. Memory card exchange has been successfully performed on 2/15.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): The ground is presently working to schedule an activity to get the orientation of the stowed sample module rotated. This is to reduce the possibility of a bubble migrating through some of the samples when they are being stowed and equilibrating for six months. The stowage procedure is being updated to reflect this desire for future operations.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Third session successfully performed on 2/14.

EuTEF (European Technology Exposure Facility): EuTEF has been successfully installed on COL External Platform during EVA-3 on 2/15 (FD9). During Week#18, 9 out of 9 instruments have been progressively checked out from ground with success. Two instruments (PLEGPAY and EXPOSE-E) have already been fully commissioned. At EuTEF platform level, troubleshooting is underway on some software issues with the DHPU MMU (Data Handling & Power Unit Mass Memory Unit), so that scripts can be loaded to operate the EuTEF platform in an automated way. The full commissioning of the remaining instruments will continue during Week #19 and the science program will begin during Week #20.

GEOFLOW: Hardware launched on 1E. The FSL (Fluid Science Laboratory) facility commissioning continues for a couple of days by Leo Eyharts, and it is currently planned to start the GEOFLOW science program during Week #20.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): In progress.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Completed 30 runs since the start of InSPACE-2. Three runs were completed on 2/19 which were performed nominally. These runs explored the low frequency response of the fluid. A progression to much larger structures with increasing field strength was observed. The final structures at the end of the last run exhibited large field distortion effects when it pulsed.

Integrated Immune: Peggy's and Leo's participation in the STS-122 (1E) Integrated Immune session is appreciated!

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Troubleshooting of the centrifuge/incubator was performed successfully on 2/15 and 2/16. The test demonstrated that the temperature can be set between +37degC and +6degC, and the KUBIK-FM3 is fully functional. Telemetry has been downlinked and is under analysis by the KUBIK Payload Developer (PD).

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: MULTIGEN-1 samples have been downloaded on STS-122 (1E) and will now be analyzed on the ground..

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): The ground appreciates Peggy's completion of the download/initialization activity off the task list. Her remaining scheduled activities include monthly downloads, two more weeks of sleep logging, and doffing the Actiwatch. Also appreciated is the additional sleep logging. 'Peggy, you continue to go above and beyond and the PI greatly appreciates it.'

SOLAR (Solar Monitoring Observatory): SOLAR has been successfully installed on COL External Platform during EVA-3 on 2/15 (FD9). Some software glitches have been repeatedly encountered with the Sun Pointing Device platform, and are currently under investigation. Some anomalies have also been encountered with a sub-system of the SOVIM instrument, which had to be power cycled. The SOLAR instruments are currently outgassing (nominal activity), before the science program is started.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): In progress.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: In progress.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): WAICO#1 hardware launched on 1E. WAICO#1 items (ATCS with chemicals, Cultivation Box Bowls with Agar) transferred to MELFI on 2/10 (FD4). Two transfer activities of 2 x 4 WAICO-#1 Seeds containers to MELFI have been performed on 2/11 to start Arabidopsis seeds vernalization phase. The BIOLAB commissioning is nearing completion. The transfer of cold stowed items from MELFI to BIOLAB was completed yesterday (2/22), and WAICO#1 experiment is currently planned to start on 02/26.

CEO (Crew Earth Observation): Ongoing.

CEO photo targets uplinked for today were Dhaka, Bangladesh (the population of greater Dhaka, the Bangladeshi capital, is 11 million, making it the 11th largest city on Earth. Images of the margins of the city are requested. The city can be hard to identify: the crew was to start looking right as ISS passed over the very wide Brahmaputra River on the India-Bangladesh border. Then they were to look on a branch river nearer track than the Brahmaputra), Ganges River Delta (fast change of river course location and island shape typify this major world delta. Requested was a mapping swath of overlapping images aimed right of track. An extreme example of fast change was the sudden appearance offshore of New Moore Island (3.5 km long) in 1970. The island appeared immediately after the storm surge and floods generated by Cyclone Bhola in November 1970. This cyclone, although only a category 3 storm, was the most destructive on record. It caused the movement of vast amounts of sediment. The island continues to grow), and East Haruj Megafans, Libya (ISS passed over the center of this remote Saharan target. A mapping swath taken near nadir was requested. The almost featureless desert landforms of these 'desert flats' have been recently recognized as vast spreads of river sediment (megafans). The sediment has been transported hundreds of km from the Tibesti Mts. far to the south. Now-dry river courses can be traced from the mountains. These rivers undoubtedly flowed during wet climate phases in the past.)

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use